### v2

@@ Abstract
Medical anomaly detection aims to identify abnormal findings using only normal training data, playing a crucial role in health screening and recognizing rare diseases. Reconstruction-based methods, particularly those utilizing autoencoders (AEs), are dominant in this field. They work under the assumption that AEs trained on only normal data cannot reconstruct unseen abnormal regions well, thereby enabling the anomaly detection based on reconstruction errors. However, this assumption does not always hold due to the mismatch between the reconstruction training objective and the anomaly detection task objective, rendering these methods theoretically unsound. This study focuses on providing a theoretical foundation for AE-based reconstruction methods in anomaly detection. By leveraging information theory, we elucidate the principles of these methods and reveal that the key to improving AE in anomaly detection lies in minimizing the information entropy of latent vectors. Experiments on four datasets with two image modalities validate the effectiveness of our theory. To the best of our knowledge, this is the first effort to theoretically clarify the principles and design philosophy of AE for anomaly detection. Code will be available upon acceptance.
@@ 3.1 An inherent property of AE
Here $W^{(1)}\in\mathbb{R}^{D\times d},W^{(2)}\in\mathbb{R}^{d\times D},d^{(1)}\in\mathbb{R}^{d},d^{(2)}\in\mathbb{R}^{D}$ are learnable weights of $g_{1}$ and $g_{2}$. The solution of Eq. 5 definitely exits. To ensure that Eq. 4 has at least one solution, the number of independent scalar equations should not exceed the number of learnable parameters, i.e., $D^{2}\leq 2Dd\Leftrightarrow d\geq\frac{D}{2}$. On the contrary, when $d<\frac{D}{2}$, the equation has no solution. Thus, if $d<\frac{D}{2}$, the AE cannot learn identical mapping. $\square$
@@ 3.2 The optimal solution for AE
Medical abnormal images typically consist of the healthy version with an addition of lesion regions, as shown in Fig 1 ($\mathbf{X}_{a}$). Therefore, all possible normal patterns could appear in abnormal images. This can be represented as
@@
(1) For a normal input $\mathbf{X}_{n}$, according to Eq. 2, the optimal AE should satisfy $\hat{\mathbf{X}}_{n}=\mathbf{X}_{n}$. Combined with Eq. 7, we have $I(\mathbf{X}_{n};\mathbf{Z})\geq I(\mathbf{X}_{n};\hat{\mathbf{X}}_{n})=I(\mathbf{X}_{n};\mathbf{X}_{n})=H(\mathbf{X}_{n})$. Using the property of mutual information, $I(\mathbf{X}_{n};\mathbf{Z})=H(\mathbf{X}_{n})+H(\mathbf{Z})-H(\mathbf{X}_{n};\mathbf{Z})\leq H(\mathbf{X}_{n})$, we have $I(\mathbf{X}_{n};\mathbf{Z})=H(\mathbf{X}_{n})$.
(2) For an abnormal input $\mathbf{X}_{a}=\mathbf{X}_{n}+\delta$, according to Eq. 3, the optimal AE should satisfy $\hat{\mathbf{X}}_{a}=\mathbf{X}_{n}$. To prevent anomaly $\delta$ from being reconstructed, $\mathbf{Z}$ should not contain abnormal information $H(\mathbf{X}_{a}|\mathbf{X}_{n})$, so $I(\mathbf{Z};\mathbf{X}_{a}|\mathbf{X}_{n})=0$. Combined with the property of mutual information, we have $I(\mathbf{Z};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{a}|\mathbf{X}_{n})+I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})$. Considering $I(\mathbf{Z};\mathbf{X}_{n})=H(\mathbf{X}_{n})$ and $I(\mathbf{X}_{a};\mathbf{X}_{n})=H(\mathbf{X}_{n})$, we simplify that $I(\mathbf{Z};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})=I(\mathbf{X}_{n};\mathbf{X}_{a})=H(\mathbf{X}_{n})$. $\square$
@@ 5 Conclusion and Discussion
Appendix 0.A Supplementary materials for experiments
0.A.1 Datasets

### v3

@@ Abstract
Medical anomaly detection aims to identify abnormal findings using only normal training data, playing a crucial role in health screening and recognizing rare diseases. Reconstruction-based methods, particularly those utilizing autoencoders (AEs), are dominant in this field. They work under the assumption that AEs trained on only normal data cannot reconstruct unseen abnormal regions well, thereby enabling the anomaly detection based on reconstruction errors. However, this assumption does not always hold due to the mismatch between the reconstruction training objective and the anomaly detection task objective, rendering these methods theoretically unsound. This study focuses on providing a theoretical foundation for AE-based reconstruction methods in anomaly detection. By leveraging information theory, we elucidate the principles of these methods and reveal that the key to improving AE in anomaly detection lies in minimizing the information entropy of latent vectors. Experiments on four datasets with two image modalities validate the effectiveness of our theory. To the best of our knowledge, this is the first effort to theoretically clarify the principles and design philosophy of AE for anomaly detection. The code is available at https://github.com/caiyu6666/AE4AD.
@@ 3.1 An inherent property of AE
Here $W^{(1)}\in\mathbb{R}^{D\times d},W^{(2)}\in\mathbb{R}^{d\times D},d^{(1)}\in\mathbb{R}^{d},d^{(2)}\in\mathbb{R}^{D}$ are learnable weights of $g_{1}$ and $g_{2}$. The solution of Eq. 5 definitely exits. If Eq. 4 has at least one solution, the number of independent scalar equations ($D^{2}$) should not exceed the number of variables ($2Dd$), i.e., $D^{2}\leq 2Dd\Leftrightarrow d\geq\frac{D}{2}$. As a contraposition, if $d<\frac{D}{2}$, Eq. 4 has no solution. Thus, if $d<\frac{D}{2}$, the AE cannot learn identical mapping. $\square$
@@ 3.2 The optimal solution for AE
Medical abnormal images typically consist of the healthy version with an addition of lesion regions, as shown in Fig. 1 ($\mathbf{X}_{a}$). Therefore, all possible normal patterns could appear in abnormal images. This can be represented as
@@
(1) For a normal input $\mathbf{X}_{n}$, according to Eq. 2, the optimal AE should satisfy $\hat{\mathbf{X}}_{n}=\mathbf{X}_{n}$. Combining Eq. 7, we have $I(\mathbf{X}_{n};\mathbf{Z})\geq I(\mathbf{X}_{n};\hat{\mathbf{X}}_{n})=I(\mathbf{X}_{n};\mathbf{X}_{n})=H(\mathbf{X}_{n})$. Using the property of mutual information and joint entropy [12], we have $I(\mathbf{X}_{n};\mathbf{Z})=H(\mathbf{X}_{n})+H(\mathbf{Z})-H(\mathbf{X}_{n},\mathbf{Z})\leq H(\mathbf{X}_{n})$. Hence, $I(\mathbf{X}_{n};\mathbf{Z})=H(\mathbf{X}_{n})$.
(2) For an abnormal input $\mathbf{X}_{a}=\mathbf{X}_{n}+\delta$, according to Eq. 3, the optimal AE should satisfy $\hat{\mathbf{X}}_{a}=\mathbf{X}_{n}$. To prevent anomaly $\delta$ from being reconstructed, $\mathbf{Z}$ should not contain abnormal information $H(\mathbf{X}_{a}|\mathbf{X}_{n})$, so $I(\mathbf{Z};\mathbf{X}_{a}|\mathbf{X}_{n})=0$. Combining the property of mutual information [12], we have $I(\mathbf{Z};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{a}|\mathbf{X}_{n})+I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})$. Considering $I(\mathbf{Z};\mathbf{X}_{n})=H(\mathbf{X}_{n})$ and $I(\mathbf{X}_{a};\mathbf{X}_{n})=H(\mathbf{X}_{n})$, we simplify that $I(\mathbf{Z};\mathbf{X}_{a})=I(\mathbf{Z};\mathbf{X}_{n};\mathbf{X}_{a})=I(\mathbf{X}_{n};\mathbf{X}_{a})=H(\mathbf{X}_{n})$. $\square$
@@ 5 Conclusion and Discussion
{credits}
5.0.1 Acknowledgements
This work was supported by the Research Grants Council of Hong Kong (No. R6003-22 and T45-401/22-N).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
@@ 5.0.1 Acknowledgements
Appendix 0.A Datasets